# EEG-MACS: Manifold Attention and Confidence Stratification for EEG-based Cross-Center Brain Disease Diagnosis under Unreliable Annotations

### v2

@@ Abstract.
Cross-center data heterogeneity and annotation unreliability significantly challenge the intelligent diagnosis of diseases using brain signals. A notable example is the EEG-based diagnosis of neurodegenerative diseases, which features subtler abnormal neural dynamics typically observed in small-group settings. To advance this area, in this work, we introduce a transferable framework employing Manifold Attention and Confidence Stratification (MACS) to diagnose neurodegenerative disorders based on EEG signals sourced from four centers with unreliable annotations. The MACS framework’s effectiveness stems from these features: 1) The Augmentor generates various EEG-represented brain variants to enrich the data space; 2) The Switcher enhances the feature space for trusted samples and reduces overfitting on incorrectly labeled samples; 3) The Encoder uses the Riemannian manifold and Euclidean metrics to capture spatiotemporal variations and dynamic synchronization in EEG; 4) The Projector, equipped with dual heads, monitors consistency across multiple brain variants and ensures diagnostic accuracy; 5) The Stratifier adaptively stratifies learned samples by confidence levels throughout the training process; 6) Forward and backpropagation in MACS are constrained by confidence stratification to stabilize the learning system amid unreliable annotations. Our subject-independent experiments, conducted on both neurocognitive and movement disorders using cross-center corpora, have demonstrated superior performance compared to existing related algorithms. This work not only improves EEG-based diagnostics for cross-center and small-setting brain diseases but also offers insights into extending MACS techniques to other data analyses, tackling data heterogeneity and annotation unreliability in multimedia and multimodal content understanding. We have released our code here: https://github.com/ICI-BCI/EEG-Disease-MACS.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681645isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Modeling methodologiesccs: Applied computing Health care information systems
@@ 1. Introduction
Figure 2. Overview of the MACS framework and subject-independent experiments across diseases and centers. The Projector has a dual-head structure for latent space representation and classification. The Stratifier categorizes samples by label confidence, constraining brain variants for contrastive learning and discriminative loss for distrusted samples. Refer to Figures 3 and 4 for Switcher and Encoder, respectively.
@@ 3.1. MACS Modules
where $\delta$ scales the values to a range of 0 to 1. $f_{W_{k}}$ represents the operation of $W_{k}$ multiplied by a matrix and then by $W_{k}^{\top}$. This can be expressed as $f_{W_{k}}(X)=W_{k}XW_{k}^{\top}$. Similarly, $f_{W_{q}}$ and $f_{W_{v}}$ operate in the same manner, where $W_{q}$, $W_{k}$, and $W_{v}\in\mathbb{R}^{d\times d_{1}}$ are learned weight matrices used for the bilinear mapping. Subsequently, the fused feature matrices =$\mathcal{F}=\{\mathcal{F}_{i}\}$ are subjected to a nonlinear activation process, involving the rectification of eigenvalues, as detailed in (Huang and Van Gool, 2017). To enable the measurement of these features $\mathcal{F}$ in Euclidean space for further processing, the embeddings $\hat{\mathcal{F}}$ are obtained by mapping $\mathcal{F}$ onto the tangent space, followed by flattening and concatenating them together.
@@
The high-level features encoded by $f_{En}$ are fed into a dual-branch projector. In this setup, one branch, $f_{Pl}$, is dedicated to reducing the feature dimension for confidence estimation and contrastive learning, obtaining $\mathcal{Z}=f_{Pl}(\hat{\mathcal{F}})$, while the other branch, $f_{Pc}$, functions as a classifier for discrimination purposes, obtaining $\mathcal{E}=f_{Pc}(\hat{\mathcal{F}})$.
@@
This module $f_{St}$ plays a crucial role in assessing the reliability of the learned representations, which in turn influences the learning mechanism of the MACS framework. It categorizes the confidence levels of samples based on $\mathcal{Z}$ through the following steps:
@@
    where $\mathcal{N}_{i}$ denotes the neighbourhood of K samples to $z_{i}$. The indicator function, denoted as $\mathds{1}_{B}$, is a function that returns $1$ when the condition $B$ is satisfied and 00 otherwise.
@@
1:$\bm{A}_{d,T}$ with unreliable annotation $Y$, maximum epochs $T_{\text{max}}$
@@
8:         Encoder: ^$\hat{\mathcal{F}}=f_{En}(A_{d,T_{s}})$
9:         Projector: $\mathcal{Z}=f_{Pl}(\hat{\mathcal{F}})$, $\mathcal{E}=f_{Pc}(\hat{\mathcal{F}})$
@@
13:         Encoder: ^$\hat{\mathcal{F}}=f_{En}(A_{d,T_{s}})$
14:         Projector: $\mathcal{Z}=f_{Pl}(\hat{\mathcal{F}})$, $\mathcal{E}=f_{Pc}(\hat{\mathcal{F}})$
@@ 3.2. MACS Constraints
MACS constraints are dependent on the confidence stratification adaptively assessed by the Stratifier module during training, which is pivotal in stabilizing representation learning in the context of unreliably annotated data. Intuitively, they regulate the behavior of the modules, influencing the forward propagation of features and the computation of backward losses, as depicted in Figure 2. Four types of constraints are implemented within the framework: they are applied to both the Augmentor and Switcher modules, and also influence the multi-view contrastive loss and discriminative loss. Those constraints work as follows:
@@ 3.3.2. Discriminative Loss
$\tilde{Y}^{(n)}=$$\begin{cases}{Y}^{(n)},&\text{if }A_{d,T_{s}}^{(n)}\in A_{crd},\\ \hat{Y}^{(n)},&\text{if }A_{d,T_{s}}^{(n)}\in A_{dst}.\end{cases}$

### v3

@@ Abstract.
Cross-center data heterogeneity and annotation unreliability significantly challenge the intelligent diagnosis of diseases using brain signals. A notable example is the EEG-based diagnosis of neurodegenerative diseases, which features subtler abnormal neural dynamics typically observed in small-group settings. To advance this area, in this work, we introduce a transferable framework employing Manifold Attention and Confidence Stratification (MACS) to diagnose neurodegenerative disorders based on EEG signals sourced from four centers with unreliable annotations. The MACS framework’s effectiveness stems from these features: 1) The Augmentor generates various EEG-represented brain variants to enrich the data space; 2) The Switcher enhances the feature space for trusted samples and reduces overfitting on incorrectly labeled samples; 3) The Encoder uses the Riemannian manifold and Euclidean metrics to capture spatiotemporal variations and dynamic synchronization in EEG; 4) The Projector, equipped with dual heads, monitors consistency across multiple brain variants and ensures diagnostic accuracy; 5) The Stratifier adaptively stratifies learned samples by confidence levels throughout the training process; 6) Forward and backpropagation in MACS are constrained by confidence stratification to stabilize the learning system amid unreliable annotations. Our subject-independent experiments, conducted on both neurocognitive and movement disorders using cross-center corpora, have demonstrated superior performance compared to existing related algorithms. This work not only improves EEG-based diagnostics for cross-center and small-setting brain diseases but also offers insights into extending MACS techniques to other data analyses, tackling data heterogeneity and annotation unreliability in multimedia and multimodal content understanding. We have released our code here: https://github.com/ICI-BCI/EEG-MACS.
@@
journalyear: 2024copyright: acmlicensedconference: Proceedings of the 32nd ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, VIC, Australiabooktitle: Proceedings of the 32nd ACM International Conference on Multimedia ($M$ ’24), October 28-November 1, 2024, Melbourne, VIC, Australiadoi: 10.1145/3664647.3681645isbn: 979-8-4007-0686-8/24/10ccs: Computing methodologies Modeling methodologiesccs: Applied computing Health care information systems
@@ 1. Introduction
Figure 2. Overview of the MACS framework and subject-independent experiments across diseases and centers. The Projector has a dual-head structure for latent space representation and classification. The Stratifier categorizes samples by label confidence, constraining brain variants for contrastive learning, and discriminative loss for distrusted samples. Refer to Figures 3 and 4 for Switcher and Encoder, respectively.
@@ 3.1. MACS Modules
where $\delta$ scales the values to a range of 0 to 1. $f_{W_{k}}$ represents the operation of $W_{k}$ multiplied by a matrix and then by $W_{k}^{\top}$. This can be expressed as $f_{W_{k}}(X)=W_{k}XW_{k}^{\top}$. Similarly, $f_{W_{q}}$ and $f_{W_{v}}$ operate in the same manner, where $W_{q}$, $W_{k}$, and $W_{v}\in\mathbb{R}^{d\times d_{1}}$ are learned weight matrices used for the bilinear mapping. Subsequently, the fused feature matrices $\mathcal{F}=\{\mathcal{F}_{i}\}$ are subjected to a nonlinear activation process, involving the rectification of eigenvalues, as detailed in (Huang and Van Gool, 2017). To enable the measurement of these features $\mathcal{F}$ in Euclidean space for further processing, the embeddings $\hat{\mathcal{F}}$ are obtained by mapping $\mathcal{F}$ onto the tangent space, followed by flattening and concatenating them together.
@@
[…]
@@
This module $f_{St}$ plays a crucial role in assessing the reliability of the learned representations, which in turn influences the learning mechanism of the MACS framework.
It categorizes the confidence levels of samples based on $\mathcal{Z}$ through the following steps:
@@
    where $\mathcal{N}_{i}$ denotes the neighbourhood of K samples to $z_{i}$. The indicator function, denoted as $\mathds{1}_{B}$, is a function that returns $1$ when the condition $B$ is satisfied and 0 otherwise.
@@
[…]
@@
8:   Encoder: $\hat{\mathcal{F}}=f_{En}(A_{d,T_{s}})$
[…]
@@
13:   Encoder: $\hat{\mathcal{F}}=f_{En}(A_{d,T_{s}})$
[…]
@@ 3.2. MACS Constraints
MACS constraints are dependent on the confidence stratification adaptively assessed by the Stratifier module during training, which is pivotal in stabilizing representation learning in the context of unreliably annotated data. Intuitively, they regulate the behavior of the modules, influencing the forward propagation of features and the computation of backward losses, as depicted in Figure 2. Four types of constraints are implemented within the framework: they are applied to both the Augmentor and Switcher modules and also influence the multi-view contrastive loss and discriminative loss. Those constraints work as follows:
@@ 3.3.2. Discriminative Loss
[…]
@@ Acknowledgements.
Appendix A Data Details
A.1. EEG Acqusition
Mild Cognitive Impairment (MCI) Dataset
The MCI dataset was acquired from a local hospital using a 64-channel Brain Products EEG system. This system was configured as Figure A1 with the following electrodes: Fp1, Fpz, Fp2, AF7, AF3, AFz, AF4, AF8, F7, F5, F3, F1, Fz, F2, F4, F6, F8, FT7, FC5, FC3, FC1, FCz, FC2, FC4, FC6, FT8, T7, C5, C3, C1, Cz, C2, C4, C6, T8, TP9, TP7, CP5, CP3, CP1, CPz, CP2, CP4, CP6, TP8, TP10, P7, P5, P3, P1, Pz, P2, P4, P6, P8, PO7, PO3, POz, PO4, PO8, O1, Oz, O2, and Iz. The electrodes at AFz and FCz were designated as ground and reference, respectively, and were excluded from the data analysis. A sampling rate of 5000 Hz was utilized during the EEG recording process. Data from the ground and reference channels were excluded from the dataset. A total of one hundred participants with cognitive decline were initially recruited. Following the inspection of medication history and other case-relevant factors (conducted by clinical physicians) as well as strict age-matching with the control group (carried out by engineering personnel), a final cohort comprising 46 subjects in the MCI group and 43 subjects in the HC group was deemed eligible for the study. Participants were instructed to sit comfortably and maintain a state of eyes-closed relaxation for an 8-minute recording session.
Alzheimer’s Disease (AD) Dataset
The AD dataset was procured from clinical monitoring at an external center, utilizing a 16-channel Symtop EEG amplifier. This system employed the International 10-20 system for electrode distribution, a configuration derived from the more comprehensive 10-10 system but with fewer channels, specifically: Fp1, Fp2, F7, F3, Fz, F4, F8, T7, C3, Cz, C4, T8, P7, P3, Pz, P4, P8, O1, and O2. The electrodes utilized for the AD data collection are marked in a darker color in Figure A1. Data capture for each participant lasted one minute, operating at a sampling frequency of 1,024 Hz. The linked earlobes, A1 and A2, served as the reference points. Under expert supervision, each participant contributed two samples—one with eyes closed and the other with eyes open. Each sample was ensured to have a minimum effective duration of 8 seconds, devoid of any significant artifacts or poor recording intervals.
Figure A1. Topographic mapping of electrode positions utilized for EEG dataset collection. The electrodes F9, F10, FT9, FT10, P9, P10, PO9, PO10, O9, and O10 were not included in the 64-channel Brain Products EEG system used for the Mild Cognitive Impairment (MCI) dataset collection. Additionally, the electrodes Fpz, F9, F10, P9, P10, PO9, PO10, O9, O10, and Iz were excluded from the 64-channel Brain Vision EEG system used for the collection of two Parkinson’s Disease (PD) datasets. The electrodes marked in darker color indicate inclusion in the 16-channel Symtop EEG system for the Alzheimer’s Disease (AD) data collection.
Parkinson’s Disease (PD) Datasets
Two PD datasets were collected from the University of New Mexico (NMU Center) and the University of Iowa (IU Center), employing a 64-channel Brain Vision EEG system. This system utilized the International 10-10 system for electrode distribution, as depicted in Figure A1, encompassing the following channels: Fp1, Fp2, AF7, AF3, AFz, AF4, AF8, F7, F5, F3, F1, Fz, F2, F4, F6, F8, FT9, FT7, FC5, FC3, FC1, FCz, FC2, FC4, FC6, FT8, FT10, T7, C5, C3, C1, Cz, C2, C4, C6, T8, TP9, TP7, CP5, CP3, CP1, CPz, CP2, CP4, CP6, TP8, TP10, P7, P5, P3, P1, Pz, P2, P4, P6, P8, PO7, PO3, POz, PO4, PO8, O1, Oz, and O2. During EEG data acquisition, the sampling rate was set to 500 Hz. Concurrently, the online reference was configured to channels CPz and Pz for the NMU and IU centers, respectively. Data from these reference channels were excluded from the publicly accessible repository. At the NMU Center, resting-state EEG recordings were obtained from 54 subjects under both eyes-open and eyes-closed conditions. At the IU Center, recordings were conducted with 28 subjects solely under the eyes-open condition. The PD and Normal Control (NC) groups are well-balanced according to the participant numbers across both datasets. To critically evaluate the adaptability of the proposed MACS framework under varying data collection conditions, it was initially trained using data from closed-eye sessions at the NMU center. Subsequently, the performance of the trained model was assessed by applying it directly to open-eye data from the IU center.
Ethical Declaration
The acquisition of EEG data was conducted in strict adherence to ethical guidelines, with approval from the relevant ethical committee, which remains confidential to comply with anonymity rules. Informed consent was duly obtained from all participants prior to their involvement in the study.
A.2. EEG Preprocessing
To ensure efficiency, the raw EEG datasets undergo preprocessing through an automated standard pipeline, as described by Ávila et al. (2023) (Gil Ávila et al., 2023). This pipeline includes filtering, outlier detection, re-referencing, and independent component analysis, designed to eliminate artifacts attributable to low-frequency drifts, high-frequency noise, head movements, cardiac activity, and eye movements. Due to the extensive manual processing and expert monitoring conducted on the AD clinical dataset, we bypassed the standard preprocessing pipeline for this preprocessed data. A total of 63 effective electrodes were employed for the two PD datasets, each utilizing a different reference electrode. For the MCI dataset, 62 effective electrodes were used, while the AD dataset employed 16 electrodes. To ensure temporal resolution consistency, all EEG recordings were uniformly downsampled to 250 Hz. For spatial alignment in the transferability test, linear interpolation was employed before inputting the data into the MACS framework to maintain consistency across the datasets.
Appendix B Implementation Details
B.1. MACS Encoder Architecture
Here, we provide additional details on the MACS Encoder’s two fundamental structures - the convolution starter $g_{str}$ and the temporal clipper $g_{clp}$ - that are omitted from the main body of the text for brevity.
Convolution Starter
The convolution starter $g_{str}$ constitutes a spatial one-dimensional convolution layer and a temporal convolution layer. Each layer is equipped with $d$ kernels, with the spatial layer utilizing kernels of size $(d,1)$ and the temporal layer employing kernels of size $(1,25)$, where $d$ denotes the number of EEG electrodes. We selected a kernel size of 25 sampling points for the EEG’s temporal layer, corresponding to the preprocessed sampling rate of 250 Hz. This kernel size enables the convolutional filter to extract features from a receptive field representing 100 ms of neural activities. The stride for both convolution operations is set to 1, ensuring a comprehensive analysis of the input signals. This architecture is designed to effectively capture generalized spatial and temporal patterns within the EEG data.
Temporal Clipper
The output generated by the convolution starter $g_{str}$ is a feature embedding denoted by $X_{d,T_{s}}$. This embedding is partitioned into $i$ non-overlapping feature clips $\{\ X_{d,t_{i}}\}\ _{i=1}^{I}$ by the temporal clipper $g_{clp}$. Consequently, the chosen value of $I$ plays a critical role in determining the temporal scale at which the dynamic functional network is analyzed in the subsequent manifold-based dynamic attention block $g_{datt}$. For instance, to observe brain synchronization conditions at a one-second temporal scale, $I$ should be configured as $T_{s}/fs$ for the embedding $X_{d,T_{s}}$. Experiments have been conducted to evaluate the effects of the selected hyperparameter, as illustrated in Table 3 of the main text. These experiments ensure that the temporal resolution is appropriately aligned with the desired analytical scale.
B.2. Evaluation Metrics
The metrics used to assess the model’s performance include accuracy, precision, recall, and the F-1 score. Building on the model trained at the fragment level, an ensemble strategy is employed to generate subject-level predictions. This approach involves averaging the probabilities at the fragment level and then determining the subject-level label using the optimal threshold, which is identified by employing Youden’s J statistic  (Youden, 1950). It is crucial to note that this strategy is consistently applied across all comparative methods to ensure fairness in the evaluation process.
Figure A2. Qualitative comparison of the MACS framework with state-of-the-art methods: t-SNE visualization of latent distribution in Parkinson’s disease (PD) data throughout training progress.
Appendix C Extended Experimental Results
C.1. Comparison Study
To guarantee fairness in the comparative analysis, all state-of-the-art (SOTA) methods implemented identical fold divisions, fragment partitioning, cross-validation approaches, and evaluation strategies.
Table A2 delineates the comprehensive results of this comparative study, showcasing full evaluation metrics. While MACS may not achieve the highest scores in every single metric, it demonstrates superior overall performance in the recognition of both MCI and PD, surpassing SOTA methods.
Qualitatively, the MACS framework was compared with three SOTA methods by visualizing the learning progress. This was achieved using t-SNE mappings to depict the embedding features of trusted samples identified by the model across various epochs. Based on their original ground truth labels, before they were reclassified as part of the unreliable annotation sets through label interchanging, the samples identified by the model as trusted were categorized as either appropriately trusted or misplaced. The main text details the model training process on the MCI dataset, showcasing the efficiency of the MACS framework. Complementing this, Figure A2 in the Appendix displays the results obtained from the PD dataset.
For Sel-CL (Li et al., 2022), Promix (Xiao et al., 2023), and our proposed MACS framework, training began at the initial epoch, and thus we visualize the progress starting from epoch 1 with intervals of 5 epochs. For the CTW method (Ma et al., 2023), which includes a warm-up phase before training, t-SNE visualization commenced after the warm-up phases, beginning from epoch 15 with the same 5-epoch intervals.
As the model trains, the desired outcome is that more samples are correctly selected, which enhances representation learning and facilitates the formation of class-specific clusters. The observations on both the MCI dataset (Figure 5 in the main text) and the PD dataset (Figure A2) have demonstrated the efficiency of the MACS framework.
Table A1. Comparative study of state-of-the-art methods for learning with well-annotated and unreliable annotated data.
[MCI] 4-Fold Cross-Validation [PD] 3-Fold Cross-Validation
Parameters Configuration Accuracy Precision Recall F1 Configuration Accuracy Precision Recall F1
0 83.10(6.93) 78.41(17.67) 92.46(8.95) 83.64(9.09) 0 85.18(3.21) 85.46(4.78) 85.19(6.41) 85.15(3.37)
200 87.60(4.46) 87.29(12.65) 86.30(5.96) 86.39(7.58) 50 83.33(0.00) 81.76(5.09) 83.01(5.58) 83.30(0.93)
300 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 100 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
400 86.31(5.55) 85.36(12.25) 85.60(13.86) 84.71(9.69) 200 85.18(3.21) 89.17(10.10) 81.48(6.41) 84.69(2.61)
Memory Length 500 85.33(4.52) 84.67(13.37) 84.21(4.80) 84.07(7.77) 300 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
2s 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 1s 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
1s 87.65(4.31) 87.12(10.98) 87.26(10.90) 86.75(7.87) 500ms 81.48(3.20) 89.17(10.10) 74.08(16.97) 79.33(6.90)
Temporal Scale 500ms 74.06(6.09) 67.24(13.56) 94.36(7.86) 77.60(8.15) 250ms 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
15 86.46(3.87) 93.03(8.13) 76.92(18.90) 82.69(11.65) 8 83.33(0.00) 85.00(4.33) 81.48(6.41) 82.97(1.07)
20 87.60(4.46) 82.68(13.28) 93.94(7.01) 87.39(7.30) 12 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
25 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 16 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
K Value 30 87.60(4.46) 82.96(12.50) 93.94(7.01) 87.61(6.87) 20 85.18(3.21) 93.33(11.55) 77.78(11.11) 83.90(3.76)
Table A2. The class-discriminative activation heatmaps generated by the MACS framework highlight localized Regions of Interest (ROIs) for detecting abnormalities in EEG signals under MCI, AD, and PD conditions. These heatmaps were averaged across all samples within each dataset, including the MCI, AD, and PD (from both the NMU and IU centers) datasets. The scatter plots illustrate the consistency between each sample and the averaged results, statistically quantified by the Spearman coefficient ($^{*}/CC^{**}$, indicating meaningful correlation) and the corresponding p-values ($^{*}/pp^{**}$, signifying significance).
[MCI] 4-Fold Cross-Validation [PD] 3-Fold Cross-Validation
Parameters Configuration Accuracy Precision Recall F1 Configuration Accuracy Precision Recall F1
0 83.10(6.93) 78.41(17.67) 92.46(8.95) 83.64(9.09) 0 85.18(3.21) 85.46(4.78) 85.19(6.41) 85.15(3.37)
200 87.60(4.46) 87.29(12.65) 86.30(5.96) 86.39(7.58) 50 83.33(0.00) 81.76(5.09) 83.01(5.58) 83.30(0.93)
300 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 100 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
400 86.31(5.55) 85.36(12.25) 85.60(13.86) 84.71(9.69) 200 85.18(3.21) 89.17(10.10) 81.48(6.41) 84.69(2.61)
Memory Length 500 85.33(4.52) 84.67(13.37) 84.21(4.80) 84.07(7.77) 300 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
2s 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 1s 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
1s 87.65(4.31) 87.12(10.98) 87.26(10.90) 86.75(7.87) 500ms 81.48(3.20) 89.17(10.10) 74.08(16.97) 79.33(6.90)
Temporal Scale 500ms 74.06(6.09) 67.24(13.56) 94.36(7.86) 77.60(8.15) 250ms 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
15 86.46(3.87) 93.03(8.13) 76.92(18.90) 82.69(11.65) 8 83.33(0.00) 85.00(4.33) 81.48(6.41) 82.97(1.07)
20 87.60(4.46) 82.68(13.28) 93.94(7.01) 87.39(7.30) 12 83.33(0.00) 89.17(10.10) 77.78(11.11) 82.19(2.11)
25 88.74(4.61) 86.15(12.39) 91.16(5.92) 88.18(7.23) 16 87.04(3.21) 93.33(11.55) 81.48(6.41) 86.40(1.90)
K Value 30 87.60(4.46) 82.96(12.50) 93.94(7.01) 87.61(6.87) 20 85.18(3.21) 93.33(11.55) 77.78(11.11) 83.90(3.76)
Figure A3. The class-discriminative activation heatmaps generated by the MACS framework highlight localized Regions of Interest (ROIs) for detecting abnormalities in EEG signals under MCI, AD, and PD conditions. These heatmaps were averaged across all samples within each dataset, including the MCI, AD, and PD (from both the NMU and IU centers) datasets. The scatter plots illustrate the consistency between each sample and the averaged results, statistically quantified by the Spearman coefficient ($^{*}/CC^{**}$, indicating meaningful correlation) and the corresponding p-values ($^{*}/pp^{**}$, signifying significance).
C.2. Hyperparameter Tuning Study
Memory Length
The multi-view contrastive loss is a pivotal component in our MACS framework, as it significantly depends on the adequacy of positive and negative pairs for effective contrastive learning. However, the prevalent scarcity of clinical data poses a substantial challenge. To circumvent the constraints imposed by batch size and to capitalize on the more extensively available data during the learning process, we adopt a storage strategy as outlined by (Ortego et al., 2021). This strategy involves retaining the preceding $M$ features in memory, with $M$ denoting the memory length. As illustrated in Table A2, the influence of memory length on the model’s performance was assessed for both MCI and PD datasets via an N-fold cross-validation experiment. This hyperparameter tuning study determined the optimal memory lengths to be 300 for the MCI dataset and 100 for the PD dataset. The results reveal a bell-shaped relationship between memory length and model performance, suggesting an optimal point appears to correlate with the dataset size, as evidenced by the MCI dataset containing approximately three times more EEG fragments than the PD dataset.
k-Nearest Neighbors
The parameter $K$ determines the number of nearest neighbors selected to infer potential labels, using the k-Nearest Neighbors algorithm in the Stratifier $f_{St}$ module of the MACS framework. A thorough exploration of $K$, employing N-fold cross-validation, is summarized in Table A2. In the experiments, the minimum value of $K$ and the interval are contingent upon the batch size of MCI and PD datasets. Similar to memory length, $K$ exhibits an ’inverted-U relationship’. Thus, $K$’s optimal configuration should be well-coordinated with the batch sizes.
Temporal Scale
As delineated in Appendix B.1, the variable $I$ dictates the temporal scale employed for analyzing characteristics within the dynamic functional networks. To this end, a series of experiments were conducted, exploring the effects of temporal scales, extending from the millisecond level to the second level. The findings presented in Table A2 indicate that adopting a second-level temporal scale is necessary for achieving robust and improved outcomes. Specifically, the temporal scale was configured to 2 seconds for the MCI dataset and 1 second for the PD dataset. The decision to configure the maximum temporal scale for PD data at 1 second was informed by the minimal length of PD data, which stands at 2 seconds. This configuration was strategically chosen to ensure that a minimum of one sample is capable of yielding two distinct data segments, thus meeting the structural prerequisites of our model. The emergence of more effective EEG markers at the larger temporal scales, as presented in Table A2, may be attributable to the ’slowing’ phenomenon in brain activity, which is more prominently manifested at broader temporal scales in patients with neurodegenerative diseases (D’Atri et al., 2021; Jones et al., 2022)
C.3. Region of Interest Localized by MACS
We applied a class-discriminative activation algorithm inspired by the Grad-CAM method (Selvaraju et al., 2017) to localize the region of interest (ROI) identified by the MACS framework while analyzing abnormalities in MCI, AD, and PD based on EEG signals.
To enhance our understanding, we computed the average activation across all samples within each dataset, resulting in the generation of four ROI heatmaps corresponding to MCI, AD, PD (MNU center), and PD (IU center). Concurrently, we aimed to assess the uniformity of the sample-specific responses relative to their respective averaged ROI activation maps. To achieve this, we utilized the Spearman rank correlation method, as detailed in the referenced pipeline  (Song et al., 2022), calculating the correlation coefficients between the averaged activation maps and each sample, along with an overall quantification metric termed $ROI_{f}$. The higher $ROI_{f}$ indicates greater consistency between all samples and their average result.
Here, a Spearman correlation coefficient ($C_{Spearman}$) of 1 (or -1) indicates a perfect positive (or negative) correlation, while a coefficient of 0 signifies the absence of correlation. In each heatmap’s right panel, the visualization includes blue triangle markers that identify individual samples. The correlation significance within these heatmaps is demarcated by two lines: a dotted green line at $C_{Spearman}=0.3$ differentiates between negligible and meaningful correlations, and a solid green line at $C_{Spearman}=0.5$ highlights correlations ranging from moderate to perfect. To establish the statistical relevance of these correlations, p-values were calculated. Correlations with p-values below 0.05 are marked by a solid red line, indicating statistical significance, while a dotted red line at p = 0.01 delineates a more rigorous threshold for significance.
Similar patterns are observed in conditions of MCI and AD, notably with significant ROIs in the parietal-occipital and frontal-temporal lobes. This may be indicative of a pathological continuum from MCI to AD, with MCI being an earlier stage in the progression towards the more pronounced dementia typical of AD. This evidence supports that the MACS framework effectively captures the core patterns in AD, potentially providing EEG evidence of compensatory mechanisms in the early stage (Gaubert et al., 2019). Analogous ROIs, such as the right frontal-temporal lobes and left temporal-parietal lobes, are observed in PD conditions, despite the data originating from two different centers. This observation suggests that the model is capable of capturing shared features across diverse datasets, thereby enhancing its transferability.
C.4. Speed and Parallelism
We assessed the latency of our framework compared to state-of-the-art (SOTA) frameworks, as shown in Table A3. Although our framework is conceptualized as a dual-branch system, the parameters are shared between modules, as illustrated in Figure 2. This dual-branch concept is realized through our contrastive losses, enabling parallelism. To ensure fair speed comparisons in weakly supervised learning, we substituted the encoders in SOTA frameworks with ours. MACS sacrifices approximately 8 ms compared to CTW and Promix to enhance performance robustness through neighbor-based confidence evaluation in the Stratifier. We doubled the speed of MACS compared to Sel-CL by implementing a selection mechanism in the Switcher that reduces pairwise samples for contrastive learning. Additionally, we compared our Encoder with the original encoders in the SOTA frameworks to demonstrate its effectiveness in reducing computational costs. Overall, our framework achieves millisecond-level inference speeds, ensuring practical application and deployment.
Table A3. Comparison of Framework Latency, Encoder Size, and Encoder Computing Complexity.
Frameworks CTW Promix Sel-CL MACS
Latency (ms) 7.39(1.42) 8.61(0.67) 37.35(1.89) 16.2(1.22)
Encoders FCN Dual ResNet ResNet Manifold Attention
Size (MB) 1.01 4.55 2.77 0.95
FLOPs (G) 0.13 0.27 1.10 0.05
Appendix D Data Availability
The PD datasets are publicly accessible and can be downloaded from http://predict.cs.unm.edu/. Access to the MCI and AD datasets will be granted upon request to the corresponding authors and subject to the approval of the collaborating hospitals.
Appendix E Code Availability
To facilitate the reproduction of our experiment, we have made the core code available at https://github.com/ICI-BCI/EEG-MACS.